OpenStudy (anonymous):

if x, y, and z are positive numbers and 2x = 3y = 4z then the value of x+y+z is how many times the value of x

OpenStudy (anonymous):

2x=3y=4z=k(say)\[x=\frac{ k }{ 2 },y=\frac{ k }{ 3 },z=\frac{ k }{ 4 }\] \[x+y+z=\frac{ k }{ 2 }+\frac{ k }{ 3 }+\frac{ k }{ 4 }=\frac{ 6k+4k+3k }{ 12 }=\frac{ 13 }{ 12 }k=\frac{ 13 }{ 12 }*2x=?\]

OpenStudy (jmark):

x + y + z = x + 2/3x + 1/2x = 13/6x Follow@TutorVista https://twitter.com/tutorvista